### v2

@@ 3. Experiments
We evaluate Oreo across five open-domain question-answering (ODQA) tasks, comparing its performance against several baselines. Our experiments holistically assess the quality of reconstructed context by Oreo from three key aspects: efficiency, effectiveness, and robustness, using a range of metrics designed to measure the performance of downstream tasks. Our primary emphasis is on short-term factual QA tasks, where the answers to queries are typically concise and consist of only a few tokens. These tasks require LMs to identify and extract precise evidence from retrieved documents while avoiding irrelevant or distracting information. Factual QA tasks are particularly sensitive to context quality, making them an ideal benchmark for evaluating the performance of Oreo . In this section, we provide details of tasks and datasets (§3.1), baselines (§3.2) and experiment setup (§3.3).
@@ 3.1. Datasets and Tasks
Single-hop question answering. We employ PopQA (PQA) (Mallen et al., 2023), NaturalQuestions (NQ) (Kwiatkowski et al., 2019), and TriviaQA (TQA) (Joshi et al., 2017) to experiment with the single-hop open-domain QA task. Each sample in these datasets has a question and annotated short extractive answers. For PopQA, we utilize the long-tail subset, which comprises 1,399 queries involving rare entities with fewer than 100 monthly Wikipedia page views (Asai et al., 2023).
Multi-hop question answering. We also test Oreo on more complex QA scenarios, specifically using multi-hop datasets HopotQA (HQA) (Yang et al., 2018) and 2WikiMultiHopQA (2WQA) (Ho et al., 2020). Each question in these datasets requires reasoning over multiple articles.
@@
Evaluation metrics. Following previous study (Wang et al., 2023a), we assess extractive QA performance (PQA, NQ, and TQA) using the Exact Match (EM) metric, while abstractive QA performance (HQA and 2WQA) is measured using unigram F1.
@@ 3.2. Baselines
For comparison, we focus on evaluating how effectively Oreo enhances vanilla RAG systems treating both the retriever and generator as black-box components, acknowledging that they may be imperfect and not allowed to be fine-tuned. We compare the performance of downstream tasks using five configurations: (1) Query only. The answer generation is performed by using only the query without incorporating any retrieved context. This mostly relies on the internal knowledge of LMs. (2) Original full content. The context for answer generation is the sequential concatenation of all retrieved document chunks. This setup uses raw, unprocessed retrieval results. (3) Passage-level filtering. In this setting, only the most relevant chunks are selected as context. Specifically, the chunk that is best-ranked and contains the ground-truth answer is chosen for each query. If no ground-truth answer is found, the top-ranked chunk, as determined by the retriever, is used instead. For single-hop tasks, only one passage is selected, while for multi-hop tasks, two passages are used. (4) Extraction and compression. We employ three information extraction and compression methods to extract informative sentences and derive summaries from retrieved documents. Specifically, we use conditional cross-mutual information (CXMI) (Fernandes et al., 2021) to train an LM to learn to filter out redundant context by following (Wang et al., 2023a). CXMI quantifies how much a sentence contribute to the correct answer. For compression, we employ Selective-Context (Li, 2023) and LLMLingua (Jiang et al., 2023) to compress the context. Selective-Context filters out unimportance context based on self-information while LLMLingua is based on perplexity. (5) reconstructed context by using Oreo .
@@ 3.3. Experiment setup
Retriever. To retrieve top-k document chunks for each query ($k=5$ unless otherwise specified), we employ a range of off-the-shelf retrievers, including Contriver (Lei et al., 2023), DPR (Karpukhin et al., 2020) and BM25 (Robertson and Walker, 1994). The choice of multiple retrievers ensures that the robustness of Oreo is tested against various retrieval mechanisms, each with different strengths and weaknesses. Additionally, we extend our experiments to include retrieval of the top-10 document chunks for 2WQA to examine whether Oreo ’s performance is sensitive to context length.
@@
Model and training. We employ T5-small (Raffel et al., 2020) as the backbone model for Oreo , though it is applicable to any encoder-decoder and auto-regressive models such as LLaMA. Oreo is implemented based on Transformer library (Wolf et al., 2020), with RL implementation built upon the open-sourced package RL4LM (Ramamurthy et al., 2023). For CML, we allow a maximum of 12 contrastive samples generated by Oreo and set beam size as 8. Unless otherwise specified, Oreo is trained for 5 epochs during SFT and 3 epochs for CML stages, with batch size 4/8/16 based on the dataset size, and using a learning rate of $55e-$ 5. Detailed parameter settings are listed in Table 2.
Inference. During inference, we limit Oreo to generate a maximum of 300 tokens for all datasets.
@@ 4. Results and Analysis
We seek to address the following questions through our experiments:
@@
    How does Oreo perform in RAG pipeline for QA tasks compared to alternative context configurations as listed in §3.2? (§3)
@@
    To what extent does Oreo reduce input token length while improving QA performance? (§4.2)
@@
    How sensitive is Oreo to noisy context and chunk order perturbations? (§6)
@@
    How does Oreo perform on unseen datasets? (§4.4)
@@ 4.1. Effectiveness Evaluation
Table 3. Average performance of Oreo on five QA tasks by using Flan-T5 and OPT-IML as downstream generators, compared with baselines.
No Retrieval Full Content Passage-level CXMI Selective-Context LLMLingua Oreo (Ours)
@@
Single-hop QA 0.125 0.2300 0.2698 0.2714 0.1696 0.1726 0.3616
Multi-hop QA 0.4416 0.334 0.5866 0.4626 0.346 0.4363 0.6542
Overall Performance. Table 3 reports the average performance across single-hop and multi-hop QA tasks using Flan-T5 and OPT-IML as downstream generators, with contexts obtained through various methods. The results demonstrate that Oreo consistently outperforms all baseline approaches, achieving the highest performance on both single-hop (EM) and multi-hop (F1) tasks. Notably, Flan-T5 generally delivers superior performance compared to OPT-IML.
Comparison against context configurations. Figure 3 presents the performance of different setups across five datasets (with Flan-T5 as downstream generator): using query-only inputs (without retrieval), original full content, passage-level filtering, Oreo without RL, and Oreo with RL. The results demonstrate that Oreo consistently outperforms the use of full content and passage-level filtering across all datasets. For single-hop QA tasks, Oreo achieves notable improvements in EM scores, with gains of 8.8%, 23.1%, and 37.5% on the PopQA, NQ, and TriviaQA datasets compared with using original full context respectively. The relatively smaller improvement on PopQA can be attributed to the nature of its queries, which involve rare entities. In the case of more complex multi-hop QA tasks, Oreo achieves F1 score improvements of 12.7% on HotpotQA and 19.8% on 2WQA. These improvements are comparatively less pronounced than those seen in single-hop tasks. This discrepancy likely stems from the increased task complexity inherent in multi-hop QA. The additional challenge of ensuring coherence in abstractive multi-hop reasoning from fragmented chunks underscores the potential for further optimization in Oreo ’s handling of such tasks. The experiments conducted on the 2WQA dataset using top-5 and top-10 retrieved document chunks demonstrate Oreo ’s flexibility in handling varying input lengths. The improved performance with the top-10 chunks arises from the increased likelihood of covering more passages that contain the ground-truth answer.
Overall, these results reveals Oreo ’s capability to capture essential information and filter out distracting content from retrieved document chunks, leading to improved performance in downstream factual QA tasks. The modest improvements achieved with reinforcement learning further emphasize its value in addressing knowledge inconsistencies between the retriever and generator.
Figure 3. Performance on five datasets by using query without retrieval, original full concatenation of chunks, passage-level filtering, context generated by Oreo without RL, and context generated by Oreo with RL. 2WQAk represents retrieving top-k documents for the 2WikiMultiHopQA dataset. The downstream generator is Flan-T5.
@@
Comparison against baselines. We compare the context quality generated by Oreo with that of selected extraction and compression methods. Table 4 summarizes the performance and token counts across five datasets by employing Flan-T5 as the downstream generator. From the table, it is evident that Oreo consistently achieves the best performance across all five datasets, with improvements ranging from 2.15% to 8.75% over the second-best methods. Notably, Oreo demonstrates more significant improvements on extractive single-hop tasks compared to multi-hop ones. This could due to the strict limitation of the output token count to 300. While this constraint is effective for single-hop tasks, where answers typically explicitly contained in the retrieved chunks, but it may lead to information loss for multi-hop QA tasks, which often require synthesizing evidence scattered across multiple chunks, demanding more lenient token caps to avoid truncation of reasoning chains. In addition to achieving superior performance, Oreo significantly reduces the context length provided to the downstream generator while maintaining or even enhancing task accuracy. This demonstrates that Oreo is effective at extracting and synthesizing the most relevant content from the retrieved context, thereby supporting more accurate generation with fewer tokens. Figure 4 illustrates the performance by using OPT-IML as the generator. Consistent with results of using Flan-T5, Oreo brings +0.0211 EM (PopQA), +0.0473 EM (NQ), + 0.0349 EM (TriviaQA), +0.0856 F1 (HotpotQA) and + 0.0495 F1 (2WQA) improvements compared with the second best baseline method.
Regarding extraction and compression baselines, the variation in performance reflects the differing strengths and limitations of these methods in addressing specific tasks. As reported in Table 4, the CXMI-guided model demonstrates performance closest to Oreo across most datasets among all baselines. In contrast, Selective-Context exhibits the lowest performance. This underperformance likely arises from its reliance on the self-information of lexical units (e.g tokens, phrases, or sentences), which overlooks the interdependence between compressed content and fails to create a cohesive context for downstream tasks. LLMLingua, which is based on perplexity, also faces challenges due to its lack of contextual modeling, as it does not explicitly consider contextual information during content selection. Furthermore, both LLMLingua and Selective-Context are not specifically designed for RAG scenarios. As a result, these methods lack query awareness, which limits their ability to tailor extracted context to the unique demands of QA tasks with RAG.
@@
Table 4. Summary of QA task performance and token counts using context derived from different methods. Flan-T5 is the generator. Bold numbers indicate the best performance among all methods, italics text denotes the second-best performance. The values in parentheses indicate the percentage improvement of the best-performing method over the second-best method. All datasets are tested with top-5 retrieved chunks.
@@
Oreo (Ours) 0.4682 (+ 8.75%) 108 0.4413 (+6.98%) 134 0.4257 (+8.35%) 130 0.6775 (+ 5.71%) 272 0.6384 (+ 2.15%) 103
Figure 4. Performance comparison against baselines by using OPT-IML as the generator. The meaning and method of each bar: Passage - passage-level filtering, CXMI - conditional cross-mutual information guided filter, Selective - Selective Context, Fulll - original full content, No - query without retrieval, Llmlingua - LLMLingua.
@@
Figure 5. Comparison of number of input tokens for generator and QA performance between different types of context.
[…]
tokenvsperf
@@ 4.2. Efficiency Evaluation
We evaluate the efficiency of Oreo from two aspects: the reduction in the input context length provided to the generator and its impact on QA task performance. Figure 5 provides an overview of the number of tokens sent to the downstream generator when using query-only inputs, full content,and context reconstructed by Oreo . Oreo achieves a remarkable 84% to 94% reduction in token counts while delivering a performance improvement ranging from 8.76% to 37.46% compared to using the original full content. The gains are particularly notable in extractive QA tasks (e.g NQ, TriviaQA) as shown from the steeper improvement in Figure 5 from right endpoints to peaks. The high compression rate and improved performance demonstrates Oreo ’s capability to effectively condense the retrieved context by preserving only the most critical evidence required for accurate answer generation. This also indicates the context reconstructed by Oreo is highly utilized by the generator.

### v3

@@ 3. Experiments
We evaluate Oreo across five open-domain question-answering (ODQA) tasks, comparing its performance against a suite of baselines. Our experiments holistically assess the quality of reconstructed context by Oreo along five critical dimensions: efficiency, effectiveness, robustness, faithfulness and completeness. Our primary emphasis is on short-term factual QA tasks, where the answers are typically concise in a few tokens. These tasks are sensitive to context quality and require precise evidence identification and summarization, making them an ideal benchmark for evaluating the performance of Oreo . In this section, we provide details of tasks and datasets (§3.1), baselines (§3.2) and experiment setup (§3.3).
@@ 3.1. Datasets and Tasks
Datasets. We evaluate Oreo on both single-hop and multi-hop open-domain question answering tasks. For single-hop QA, we use PopQA (PQA) (Mallen et al., 2023), NaturalQuestions (NQ) (Kwiatkowski et al., 2019), and TriviaQA (TQA) (Joshi et al., 2017). For multi-hop QA, we test Oreo on the more complex HopotQA (HQA) (Yang et al., 2018) and 2WikiMultiHopQA (2WQA) (Ho et al., 2020), where each question requires reasoning over multiple articles.
@@
Evaluation metrics. Following previous studies, eg.(Wang et al., 2023a), we assess extractive QA performance (PopQA, NQ, and TriviaQA) using the Exact Match (EM) metric, while abstractive QA performance (HotpotQA and 2WQA) are measured using unigram F1.
We provide detailed statistics and experimental setups for each dataset in Appendix B.
@@ 3.2. Baselines
For comparison, we focus on evaluating how effectively Oreo enhances vanilla RAG systems treating both the retriever and generator as black-box components, acknowledging that they may be imperfect and not allowed to be fine-tuned. We compare the performance of downstream tasks using five configurations:
  1. (1)
    Query only. The answer generation is performed by using only the query without incorporating any retrieved context. This mostly relies on the internal knowledge of LMs
  2. (2)
    Original full content. The context for answer generation is the sequential concatenation of all retrieved document chunks. This setup uses raw, unprocessed retrieval results
  3. (3)
    Passage-level filtering. Only the most relevant chunks are selected as context. Specifically, the chunk that is best-ranked is chosen for each query. For single-hop tasks, only one passage is selected, while for multi-hop tasks, two passages are used
  4. (4)
    Extraction and compression. We employ eight state-of-the-art information extraction and compression methods to select informative sentences and generate concise summaries from retrieved documents. Specifically:
    1. (a)
      CXMI-trained model, following (Wang et al., 2023a), uses conditional cross-mutual information (CXMI) (Fernandes et al., 2021) to train a language model to filter redundant context by quantifying each sentence’s contribution to the correct answer
    2. (b)
      Selective-Context (Li, 2023) removes uninformative content based on self-information
    3. (c)
      LLMLingua (Jiang et al., 2023) and LLMLingua-2 (Pan et al., 2024) apply perplexity-based compression to retain sentences that most enhance answer likelihood
    4. (d)
      xRAG (Cheng et al., 2024) encodes passages into a single embedding token, integrating them via modality fusion into the LM’s representation space
    5. (e)
      CompAct (Yoon et al., 2024) is a progressive compression framework that preserves query-aware content
    6. (f)
      EXIT (Hwang et al., 2024) employs an adaptive extractive pipeline to select context based on query relevance
    7. (g)
      Refiner (Li et al., 2024) uses a decoder-only LLM to extract verbatim query-relevant spans, organizing them by interdependencies
  5. (5)
    reconstructed context by using Oreo
@@ 3.3. Experiment setup
Retriever. To retrieve top-k document chunks for each query ($k=5$ unless otherwise specified), we employ a range of off-the-shelf retrievers, including Contriver (Lei et al., 2023), DPR (Karpukhin et al., 2020) and BM25 (Robertson and Walker, 1994). The choice of multiple retrievers ensures that the robustness of Oreo is tested against various retrieval mechanisms, each with different strengths and weaknesses. Additionally, we extend our experiments to include retrieval of the top-10 document chunks for 2WikiMultihopQA to examine whether Oreo ’s performance is sensitive to context length.
@@
Model and training. We employ T5-small (Raffel et al., 2020) as the backbone model for Oreo , though it is applicable to any encoder-decoder and auto-regressive models such as LLaMA. Oreo is implemented based on Transformer library (Wolf et al., 2020), with RL implementation built upon the open-sourced package RL4LM (Ramamurthy et al., 2023). For CML, we allow a maximum of 12 contrastive samples generated by Oreo and set beam size as 8. Unless otherwise specified, Oreo is trained for 5 epochs during SFT and 3 epochs for CML stages, with batch size 4/8/16 based on the dataset size, and using a learning rate of $55e-$ 5. Detailed parameter settings are listed in Appendix C.
Inference. During inference, we perform ablation studies by varying the number of tokens produced by Oreo to assess its impact on performance.
@@ 4. Results and Analysis
We seek to answer the following questions through experiments:
@@
    How does Oreo perform in the RAG pipeline for QA tasks compared to alternative context configurations outlined in §3.2? (§2)
@@
    To what extent does Oreo reduce input token length and balance inference latency while improving QA performance? (§4.2)
@@
    How robust is Oreo to noisy contexts and perturbations in chunk order? (§5)
@@
    How well does Oreo generalize to out-of-distribution datasets unseen during training? (§4.4)
  5. (5)
    How effective is Oreo in generating context that are complete to the query and faithful to the retrieved chunks? (§4.5)
  6. (6)
    How does the number of tokens in the reconstructed context affect downstream QA performance? (§4.6)
@@ 4.1. Effectiveness Evaluation
Table 2. Average performance of Oreo across five QA benchmarks using Flan-T5 and OPT-IML as downstream generators, compared with baseline methods. SC denotes Selective-Context. Performance on single-hop and multi-hop QA tasks is evaluated using Exact Match and Unigram F1, respectively. Bold values indicate the best results.
No Retrieval Full Passage CXMI SC LLMLingua LLMLingua-2 xRAG CompAct EXIT Refiner Oreo (Ours)
@@
Single-hop QA 0.125 0.2300 0.2698 0.2714 0.1696 0.1726 0.2461 0.2297 0.3142 0.3075 0.3160 0.3616
Multi-hop QA 0.4416 0.334 0.5866 0.4626 0.346 0.4363 0.5501 0.4818 0.5865 0.5828 0.5834 0.6542
Overall Performance. Table 2 reports the average performance across single-hop and multi-hop QA tasks using Flan-T5 and OPT-IML as downstream generators, with contexts obtained through various retrieval, extraction, and compression strategies. Across all settings, Oreo consistently outperforms other approaches, achieving the highest performance on both single-hop (Exact Match) and multi-hop (F1) tasks. Flan-T5 generally delivers superior performance compared to OPT-IML, likely due to its more advanced instruction tuning.
Comparison against context configurations. Figure 3 presents the performance of different setups across five datasets (with Flan-T5 as downstream generator): using query-only inputs (without retrieval), original full content, passage-level filtering, Oreo with and without RL. The results demonstrate that Oreo surpasses all other configurations across five datasets using Flan-T5. For single-hop QA tasks, Oreo achieves notable improvements in EM scores, with gains of 8.8%, 23.1%, and 37.5% on the PopQA, NQ, and TriviaQA datasets compared with using original full context respectively. The relatively smaller improvement on PopQA can be attributed to the nature of its queries, which involve rare and long-tail entities. In the case of more complex multi-hop QA tasks, Oreo achieves F1 score improvements of 12.7% on HotpotQA and 19.8% on 2WQA. These improvements are comparatively less pronounced than those seen in single-hop tasks. This discrepancy likely stems from the increased task complexity inherent in multi-hop QA. The additional challenge of ensuring coherence in abstractive multi-hop reasoning from fragmented chunks underscores the potential for further optimization in Oreo ’s handling of such tasks. The experiments conducted on the 2WQA dataset using top-5 and top-10 retrieved document chunks demonstrate Oreo ’s flexibility in handling different input lengths. The improved performance with the top-10 chunks arises from the increased likelihood of covering more passages that contain the ground-truth answer.
Overall, these results reveals Oreo ’s capability to capture essential information and filter out distracting content from retrieved document chunks, leading to improved performance in downstream factual QA tasks. The modest improvements achieved with RL further emphasize its value in addressing knowledge inconsistencies between the retriever and generator.
Figure 3. Performance on five datasets by using query without retrieval, original full concatenation of chunks, passage-level filtering, context generated by Oreo with and without RL. 2WQAk represents retrieving top-k documents for the 2WQA dataset. The downstream generator is Flan-T5. Performance of PopQA, NQ and TriviaQA are measured by Exact Match and HotpotQA and 2WQA are measured by unigram F1.
@@
Comparison against baselines. We compare the quality of generated context by Oreo against a suite of representative extraction and compression methods across five QA datasets. Table 3 summarizes the performance and token counts across five datasets by employing Flan-T5 as the downstream generator. From the table, it is evident that Oreo outperforms almost all selected methods across five datasets, with improvements ranging from 0.35% to 8.58% over the second-best methods. These gains are particularly pronounced in extractive, single-hop tasks such as NQ and PopQA, where concise yet precise evidence retrieval is paramount. On multi-hop tasks, Oreo remains competitive but shows relatively smaller gains, as these tasks require complex evidence chaining and reasoning to synthesize evidence scattered across multiple chunks. In addition to achieving superior performance, Oreo significantly reduces the context length provided to the downstream generator while maintaining or even enhancing task accuracy. We further validate these findings by evaluating with OPT-IML as the downstream generator (see Fig. 4). Consistent with results of using Flan-T5, Oreo leads to the highest performance across four datasets (except HotpotQA), bringing +0.0211 EM (PopQA), +0.0405 EM (NQ), +0.0019 EM (TriviaQA), -0.0142 F1 (HotpotQA) and +0.0495 F1 (2WQA) improvements compared with the second-best baseline method.
SOTA methods observations. Among selected SOTA methods, CompAct (Yoon et al., 2024) marginally outperforms Oreo on HotpotQA, achieving a 2.26% higher F1 score. This advantage is attributed to CompAct’s incremental and iterative compression strategy, which proves beneficial for tasks requiring deep multi-hop reasoning. However, its inference latency is nearly 3X that of Oreo , presenting a trade-off between accuracy and efficiency. Other strong performers include CXMI-guided model, Refiner and EXIT, which use query-aware or contrastive objectives to maintain relevance. In contrast, Selective-Context yields the weakest results overall. This underperformance likely stems from its reliance on self-information of lexical units (eg., tokens, phrases, or sentences), which fail to capture dependencies among semantic units. Similarly, LLMLingua and LLMLingua-2, which employ perplexity-based filtering, also struggle across datasets. Their reliance on self-information and perplexity metrics, without explicit query conditioning, limits their ability to extract context tightly aligned with user queries.
@@
Table 3. Summary of QA task performance and token counts using context derived from different methods. Flan-T5 is the generator. Performance on PopQA, NaturalQuestions, and TriviaQA is evaluated using Exact Match, while HotpotQA and 2WikiMultihopQA are assessed using Unigram F1. Bold values indicate the best performance among all methods, italics text denotes the second-best performance. The values in parentheses indicate the percentage improvement of the best-performing method over the second-best method. All datasets are tested with top-5 retrieved chunks and all retrieved passages are set the same for different methods.
@@
LLMLingua-2 0.2603 252 0.1892 247 0.3573 265 0.6186 269 0.4965 279
xRAG 0.2117 31 0.2580 40 0.2893 32 0.4777 48 0.4879 36
CompAct 0.3917 142 0.265 173 0.4242 180 0.6932 174 0.5015 173
EXIT 0.3972 124 0.2301 211 0.4188 208 0.6742 180 0.5104 123
Refiner 0.4312 91 0.2677 148 0.4051 148 0.6706 131 0.5144 102
Oreo (Ours) 0.4682 (+ 8.58%) 108 0.4413 (+6.98%) 134 0.4257 (+0.35%) 130 0.6775 (-2.26%) 272 0.6384 (+ 4.55%) 103
Figure 4. Performance comparison with 95% confidence intervals against baselines using OPT-IML as the generator. Specifically, Passage denotes passage-level filtering, CXMI refers to filtering guided by conditional cross-mutual information, and Full represents the use of original content without any filtering. PopQA, NQ, and TriviaQA are evaluated with Exact Match scores, while HotpotQA and 2WQA use Unigram F1 for accuracy measurement
@@ 4.2. Efficiency Evaluation
We assess the efficiency of Oreo from two key perspectives: (1) the trade-off between context length reduction and downstream QA performance, and (2) the trade-off between end-to-end inference latency and QA performance.
Figure 5 illustrates the number of tokens forwarded to the downstream generator and the total inference latency, which includes both Oreo ’s context reconstruction and the subsequent generation time. We compare three input configurations: query-only (no retrieval), full document content, and the context reconstructed by Oreo . Oreo achieves a substantial reduction in input length, compressing the context by 84% to 94% compared to full document content. This compression is accompanied by a latency reduction of 22.98% to 43.01%, while simultaneously delivering significant performance improvements ranging from 8.76% to 37.46%. These gains are especially pronounced in extractive QA tasks (eg., NQ, TriviaQA) as shown from the steeper improvement in Figure 5 from right endpoints to peaks. The high compression rate and improved performance demonstrates Oreo ’s capability to effectively condense the retrieved context by preserving only the most critical evidence required for accurate answer generation. This also indicates the context reconstructed by Oreo is highly utilized by the downstream generator. The favorable trade-off between latency and performance underscores Oreo ’s potential for real-world applications, offering both computational efficiency and improved task accuracy for scalable, high-throughput RAG systems.